jigglypuff314 (jigglypuff314):

\[If~\int\limits x^2cosx~dx = f(x) - \int\limits 2xsinx~dx\]then f(x) = a) 2sinx + 2xcosx b) x^2sinx c) 2xcosx - x^2sinx d) 4cosx - 2xsixs e) (2-x^2)cosx - 4sinx

jigglypuff314 (jigglypuff314):

.... >,< I think it's B from integrating the first part of it... but I don't know why or how :/

OpenStudy (ipwnbunnies):

Hint: Do you recognize how the second part can be the second part of the integrations by parts formula? :3

OpenStudy (vishweshshrimali5):

@jigglypuff314 Actually try figuring out the difference between integration by parts formula and the given integral.

jigglypuff314 (jigglypuff314):

ah so it is B? :D lel I was trying to do Tabular the first time >,<

jigglypuff314 (jigglypuff314):

u = x^2 v = sinx du = 2x dx dv = cosx dx uv - ∫v du x^2sinx - ∫2xsinx dx Thank you! :)

jigglypuff314 (jigglypuff314):

O-O whyyy integral of cosx = sinx there would be no canceling negatives... ?

OpenStudy (ipwnbunnies):

Mhm, dv = cos x dx. v = sin x + C The minus sign is correct.

OpenStudy (vishweshshrimali5):

Sorry it was correct.

OpenStudy (vishweshshrimali5):

I misread the "v"

jigglypuff314 (jigglypuff314):

hehehe xD I sometimes do that too :P Thanks for your help <3